OpenStudy (anonymous):

I WILL FAN!!!!!! Solve for x. 3 • 4 – 3 • 7 = 3(x – 7) A. –10 B. 6 C. 4 D. –3

OpenStudy (anonymous):

3 x 4 = ?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

3 x 7 = ?

OpenStudy (anonymous):

21

OpenStudy (anonymous):

12 - 21 = ?

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

I forgot the rest..Haha...

OpenStudy (anonymous):

@Vocaloid

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

Add or subtract -9 and 7?

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

I think you add -9 and -7.

Vocaloid (vocaloid):

picking up where theinspector left off:

OpenStudy (anonymous):

Haha

Vocaloid (vocaloid):

3 • 4 – 3 • 7 = -9 so -9 = 3(x-7) divide both sides by 3

Vocaloid (vocaloid):

good, so -3 = x - 7 now add 7 to both sides

OpenStudy (anonymous):

so is 4 the answer @Vocaloid

Vocaloid (vocaloid):

yeah